# Multi-Agent Diagnostics for Robustness via Illuminated Diversity

### v1

@@ Abstract.
Multi-Agent Learning, Open-Endedness, Generalisation
@@ 1. Introduction
In recent times, multi-agent systems, particularly those designed to interact with humans, have emerged as a primary model for AI deployment in real-world scenarios OpenAI (2023); Anthropic (2023); Badue et al. (2019); Touvron et al. (2023). Although there have been significant successes in simulated environments, as evidenced by deep reinforcement learning (RL) in complex multi-agent games (Silver et al., 2016; Schrittwieser et al., 2020; Vinyals et al., 2019; Berner et al., 2019; Wurman et al., 2022), the transfer from simulation to reality (sim2real) continues to pose challenges (Höfer et al., 2021; Zhao et al., 2020). Specifically, while these models demonstrate proficiency in known environments, they become highly susceptible to faulty behaviors in unfamiliar settings and adversarial situations (Samvelyan et al., 2023). Given their critical roles in human-centric applications, understanding and mitigating these susceptibilities becomes paramount for fostering more effective and reliable deployment of multi-agent AI systems in the future.
@@
Figure 1. Overview of MADRID. Operating on a discretised grid with an added dimension for reference policies, MADRID archives environment variations (or levels) characterized by representative features, e.g., $(x,y)$ coordinates of the ball position in football. During each iteration, MADRID mutates a selected level, computes regret using its associated reference policy, and reincorporates levels with higher regret into the archive, effectively generating diverse collection of adversarial levels.
@@
The Achilles’ heel of these multi-agent systems, contributing to their lack of robustness, is often their overfitting to the specific settings encountered during training (Lanctot et al., 2017). This overfitting becomes notably evident in two-team zero-sum settings where both cooperative and competitive dynamics intertwine. A primary manifestation of the overfitting between cooperative agents, especially when all agents in the group share the same set of network parameters (i.e., parameter sharing (Foerster et al., 2016)), is in the agents becoming too accustomed to their training environments, leading to a detailed coordination tailored to these specific conditions. As a consequence, when introduced to unfamiliar settings, their performance tends to falter. Concurrently, there is also an overfitting to specific opponent teams they have trained against. Instead of developing a flexible strategy that can withstand a variety of opponents, their strategies might be overly optimised to counteract the strategies of familiar adversaries. These dual forms of overfitting—both to the environment and to opponents—render such settings as perfect platforms to probe for vulnerabilities Tuyls et al. (2021). Furthermore, it is crucial to pinpoint a diverse set of adversarial scenarios for a holistic diagnostic of robustness, shedding light on possible shortcomings from various perspectives.
@@
Given these challenges, we introduce Multi-Agent Diagnostics for Robustness via Illuminated Diversity (MADRID), a novel method for systematically generating a diverse collection of adversarial settings where pre-trained multi-agent policies make strategic mistakes. To this end, MADRID employs approaches from quality-diversity (QD) (Lehman and Stanley, 2011; Cully and Demiris, 2018), a family of evolutionary algorithm that aim to generate a large collection of high-performing solutions each with their own unique characteristics.
@@
MADRID incorporates MAP-Elites (Mouret and Clune, 2015), a simple and effective QD approach, to systematically explore the vast space of adversarial settings. By discretising the search space, MADRID iteratively performs selection, mutation, and evaluations steps, endlessly refining and expanding the repertoire of high-performing adversarial scenarios within its archive (see Figure 1). A crucial feature of MADRID is its employment of the target policy’s regret— the gap in performance between the optimal and target policy—to quantify the quality of adversarial settings. Regret is shown to be an effective metric for identifying situations where RL agents underperform in both single-agent (Dennis et al., 2020; Jiang et al., 2021; Parker-Holder et al., 2022; Mediratta et al., 2023) and multi-agent (Samvelyan et al., 2023) domains. MADRID estimates a lower-bound on the true regret by utilising a collection of reference policies Vinyals et al. (2019); Garnelo et al. (2021), which are not necessarily required to be high-performing. MADRID identifies situations where these reference policies surpass the target one, thereby providing a clear illustration of superior performance in given situations.
@@
To evaluate MADRID, we concentrate specifically on one of the most challenging multi-agent domains, namely the fully decentralised 11 vs 11 variation of Google Research Football (GRF, Kurach et al., 2020). This simulated environment is based on the popular real-world sport of football (a.k.a. soccer) and requires two teams of agents to combine short-term control techniques with coordinated, long-term global strategies. GRF represents a unique combination of characteristics not present in other RL environments (Lin et al., 2023), namely multi-agent cooperation (within each team), competition (between the two teams), sparse rewards, large action and observation spaces, and stochastic dynamics. While many of the individual challenges in GRF, including multi-agent coordination (Rashid et al., 2018; Yu et al., 2022), long-term planning (Ecoffet et al., 2020) and non-transitivity (Balduzzi et al., 2019; Czarnecki et al., 2020), have been studied extensively in isolation, learning highly-competitive GRF policies has long remained outside the reach of RL methods. TiZero (Lin et al., 2023), a recent multi-agent RL approach, learned to ”master” the fully decentralised variation of GRF from scratch for the first time, using a hand-crafted curriculum, reward shaping, and self-play. Experimentally, TiZero has shown impressive results and outperformed previous methods by a large margin after an expensive training lasting 45 days on a large-scale distributed training infrastructure.
@@
We apply MADRID on GRF by targeting TiZero to diagnose a broad set of scenarios in which it commits tactical mistakes. Our extensive evaluations reveal diverse settings where TiZero exhibits a poor performance, where weaker policies can outperform it. Specifically, MADRID discovers instances where TiZero is ineffective at near the opponent’s goal, demonstrates a marked inability to comprehend the offside rule effectively, and even encounters situations of scoring accidental own goals. These findings highlight the latent vulnerabilities within even highly trained models and demonstrate that there is much room for improving the their robustness. Our analysis showcases the value of identifying such adversarial settings in offering new insights into the hidden weaknesses of pretrained policies that may otherwise appear undefeatable.
@@ 2. Background
In this work, we consider Underspecified Stochastic Games (USG), i.e., stochastic games (Shapley, 1953) with underspecified parameters of the environment. An USG game for $N$ agent environment is defined by a set of states $\mathcal{S}$, actions $\mathcal{A}_{1},...,\mathcal{A}_{N}$ and a set of observations $\mathcal{O}_{1},...,\mathcal{O}_{N}$ for each of the agents. Each agent $i$ select actions using a stochastic policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\mapsto[0,1]$. $\Theta$ defines the free parameters of the environments which are incorporated into the transition function $\mathcal{T}:\mathcal{S}\times\Theta\times\mathcal{A}_{1}\times...\times\mathcal{A}_{N}\mapsto\mathcal{S}$ which produces the next state based on the actions of all agents. Each agent $i$ receives observations $\mathbf{o}_{i}:\mathcal{S}\mapsto\mathcal{O}_{i}$ correlated with the current state and reward $r_{i}:\mathcal{S}\times\mathcal{A}_{i}\mapsto\mathbb{R}$ as a function of the state and agent’s action. The goal of each agent $i$ is to maximise its own total expected return $R_{i}=\sum_{t=0}^{T}\gamma^{t}r^{t}_{i}$ for the time horizon $T$, where $\gamma$ is a discount factor.
Each configuration of the free parameter $\theta\in\Theta$, which is often called a level (Jiang et al., 2021; Parker-Holder et al., 2022), defines a specific instantiation of the environment $\mathcal{M}_{\theta}$. For example, this can correspond to different positions of the walls in a maze, or locations of players and the ball in a football game. USG is a multi-agent variation of Underspecified POMDPs (Dennis et al., 2020) and fully observable variant of UPOSGs (Samvelyan et al., 2023).
@@
Quality-diversity (QD) is a family of methods used to find a diverse collection of solutions that are performant and span a meaningful spectrum of solution characteristics (Lehman and Stanley, 2011; Cully and Demiris, 2018). The performance of solution $x\in\mathcal{X}$ is measure using the fitness $:\mathcal{X}\mapsto\mathbb{R}$ function. The diversity of solutions is typically measured using the feature_descriptor $:\mathcal{X}\mapsto\mathbb{B}$ function that maps a solution into the feature space $\mathbb{B}=\mathbb{R}^{K}$ that describes specific characteristics of the solution, such as behavioral properties or visual appearance.
Initialise: $N$-dimensional grids for solutions $X$ and performances $\mathcal{P}$
Initialise: $n$ cells of $X$ with random solutions and corresponding cells of $P$ with their fitness
@@
MAP-Elites is a simple and effective QD method (Mouret and Clune, 2015). Here, the descriptor space $\mathbb{B}$ is discretised and represented as an initially empty $N<K$ dimensional grid (archive). The algorithm starts by generating an arbitrary collection of candidate solutions. In each iteration, a solution is randomly selected among those in the grid. A new solution is obtained by mutating the selected solution, which is then evaluated and mapped to a cell of the grid based on its feature descriptor. The solution is then placed in the corresponding cell of the grid if it has a higher fitness than the current occupant, or if the cell if it is empty. This cycle of selection, mutation, and evaluation is repeated, progressively enhancing both the diversity (coverage) and the quality (fitness) of the collection. The pseudo-code of MAP-Elites is presented in Algorithm 1.
@@ 3. MADRID
In this section, we describe Multi-Agent Diagnostics for Robustness via Illuminated Diversity (MADRID), a novel method for automatically generating diverse adversarial settings for a target pre-trained policy $\pi_{T}$. These are settings that either deceive the policy, forcing it to produce incorrect behaviour, or where the policy inherently performs poorly, deviating from the optimal behaviour. For USGs, these settings correspond to particular environment levels $\theta\in\Theta$ that have been procedurally generated.
For quantifying adversarial levels, we make use the target policy’s regret in level $\theta$, i.e., the difference in utility between the optimal policy $\pi^{*}$ and $\pi_{T}$ :
@@
where $V_{\theta}(\pi_{A},\pi_{B})=\mathop{\mathbb{E}}[\sum_{t=0}^{T}\gamma^{t}r_{t}^{A}]$ is the value of a policy $\pi_{A}$ against policy $\pi_{B}$ in $\theta$ [^2] Note that here, for the simplicity of the notation, we assume a two-team zero sum setting. $\pi_{T}$ and $\pi_{R}$ describe the policies for groups of agents, either through a centralised controller or decentralised policies that employ parameter sharing. However, MADRID can be applied for more general multi-agent settings.
Regret is a suitable metric for evaluating adversarial examples in pre-trained models. It provides a measure that directly quantifies the suboptimality of a model’s decisions. While a high regret value serves as a glaring indicator of how far off a model’s behavior is from the optimal choice, a low regret indicates the model’s decisions are closely aligned with the optimal choice. The importance of regret becomes even more pronounced when considering the varied scenarios in which a model might be deployed. Therefore, by investigating regret across diverse situations, we can not only pinpoint specific vulnerabilities of a model but also ensure the robustness in previously unseen scenarios.
@@
Input: Target policy $\pi_{T}$
Input: A collection of reference policies $\Pi_{R}$
Input: $level\_descriptor:\Theta\mapsto\mathbb{R}^{N}$ function
@@
Initialise: $N+1$-dimensional grids for levels $X$ and regret estimates $\mathcal{P}$
Initialise: $n$ cells of $X$ with randomly generated levels and corresponding estimated regret in $\mathcal{P}$
@@
MADRID casts the task of generating a diverse array of adversarial levels for each reference policy as a QD search problem. Specifically, MADRID uses MAP-Elites to systematically generate levels from $\Theta$ by discretising the feature space of levels into an $N$-dimensional grid, with an additional dimension representing the corresponding reference policy from $\Pi_{T}$. Using a discretised grid of MAP-Elites provides interpretability to the adversarial examples found in MADRID given that each cell defines specific environment parameters, alongside a reference policy which outperforms the target under these parameters.
MADRID starts by populating the grid with initial levels for each reference policy. During the iterative process, levels are selected from the grid to undergo mutation, followed by regret estimation. Each mutated level is then mapped to a specific cell in the grid based on its features and replaces the existing occupant if the mutated level has higher regret or the corresponding cell is unoccupied. This procedure ensures a thorough exploration and exploitation of the environment design space, allowing MADRID to generate levels that are both diverse and high-regret. Figure 1 illustrates this process. Algorithm 2 provides the pseudocode of the method.
@@ 4. Experimental Setting
Our experiments seek to (1) showcase the effectiveness of MADRID in generating diverse adversarial settings for a target state-of-the-art pre-trained RL model, (2) analyse the adversarial settings generated by MADRID to find key weaknesses of the target model, (3) validate the design choices of MADRID by comparing it to two ablated baselines. To this end, we evaluate MADRID on Google Research Football (GRF, Kurach et al., 2020). Given its strong performance and usage in related works, Covariance Matrix Adaptation MAP-Elites (CMA-ME, Fontaine et al., 2020) serves as the base MAP-Elites method in our experiments. We provide full environment descriptions in Appendix B and implementation details in Appendix C.
@@
We compare MADRID against two baselines: The targeted baseline uses a MAP-Elites archive but randomly samples levels from scratch, rather then evolving previously discovered high-regret levels from the grid. Consequently, it does not leverage the stepping stones to the optimisation problem (Lehman and Stanley, 2011). The random baseline samples levels randomly from scratch without maintaining an archive of high-regret levels.
@@
GRF is a complex open-source RL environment designed for training and evaluating agents to master the intricate dynamics of football, one of the world’s most celebrated sports. It offers a physics-based 3D simulation that tasks the RL policy with controlling a team of players to penetrate the opponent’s defense, while passing the ball among teammates, in order to score goals. GRF is a two-team zero-sum environment that has long been considered one of the most complex multi-agent RL benchmarks due to a unique combination of challenges (Huang et al., 2021; Wen et al., 2022; Lin et al., 2023), such as multi-agent cooperation, multi-agent competition, sparse rewards, large action and observation spaces, and stochastic dynamics. [^3] Highlighting the stochasticity of the GRF environment, a shot from the top of the box can lead to various outcomes, underscoring that not every action results in a predictable outcome.
In this work, we focus on the fully decentralised 11 vs 11 version of the environment where each of the $10$ RL agents on both sides controls an individual player on the field. [^4] The goalkeepers are controlled by the game AI. Following (Lin et al., 2023), each agents receives as observation a $268$-dimensional feature vector include own player information, player IDs, as well as information about the ball, player of the own and opponents teams, as well as general match details. The action space of agents consists of $19$ discrete actions, such as moving in 8 direction, sprinting, passing, shooting, etc.
To apply MADRID on GRF, we utilise procedurally generated levels each represented as a vector consisting of $(x,y)$ coordinates of $20$ players [^5] The goalkeepers position positions are always near their own goals. and the ball. The position of the ball on the field serves as a convenient descriptor for levels in GRF because it accommodates diverse scenarios, ranging from attacking to defending on both field halves. Therefore, we use the $x$ and $y$ coordinates of the ball as the first two environment features in MADRID. This leads to a categorisation of levels into $160$ uniformly spaced cells across the football field, as illustrated in Figure 3. Given that we are interested in evaluating TiZero in specific adversarial levels, in our experiments we restrict the episode length to $128$ steps taking place in the beginning of the game.
The third axis for the MAP-Elites archive indexes the reference policies $\Pi_{R}$. In our experiments, we make use of $48$ checkpoints of TiZero saved throughout its training (Lin et al., 2023), as well as three built-in bots in GRF with varying difficulties (easy, medium, hard). For each reference policy, we initialise the grid with randomly sampled levels that assign random locations to players and the ball. Figure 3 illustrates some of the randomly generated levels.
@@
which corresponds to the difference of cross-play and self-play values between the reference and target policies. The performance on a given level $\theta$ between two policies $\pi_{A}$ and $\pi_{B}$ is the reward for scoring a goal:
@@
Figure 4. The estimated regret and goal score rate against TiZero in Google Research Football. Illustrated throughout each iteration for 51 reference agents (a) and (c), as well as final values in (b) and (d). Standard error over 3 random seeds is shown.

### v2

@@ Abstract.
Multi-Agent Learning; Open-Endedness; Robustness
@@ 1. Introduction
In recent times, multi-agent systems, particularly those designed to interact with humans, have emerged as a primary model for AI deployment in real-world scenarios OpenAI (2023); Anthropic (2023); Badue et al. (2019); Touvron et al. (2023). Although there have been significant successes in simulated environments, as evidenced by deep reinforcement learning (RL) in complex multi-agent games (Silver et al., 2016; Schrittwieser et al., 2020; Vinyals et al., 2019; Berner et al., 2019; Wurman et al., 2022), the transfer from simulation to reality (sim2real) continues to pose challenges (Höfer et al., 2021; Zhao et al., 2020). Specifically, while these models demonstrate proficiency in known environments, they become highly susceptible to faulty behaviours in unfamiliar settings and adversarial situations (Samvelyan et al., 2023). Given their critical roles in human-centric applications, understanding and mitigating these susceptibilities becomes paramount for fostering more effective and reliable deployment of multi-agent AI systems in the future.
@@
Figure 1. Overview of MADRID. Operating on a discretised grid with an added dimension for reference policies, MADRID archives environment variations (or levels) characterized by representative features, e.g., $(x,y)$ coordinates of the ball position in football. During each iteration, MADRID mutates a selected level, computes regret using its associated reference policy, and reincorporates levels with higher regret into the archive, effectively generating a diverse collection of adversarial levels.
@@
The Achilles’ heel of these multi-agent systems, contributing to their lack of robustness, is often their overfitting to the specific settings encountered during training (Lanctot et al., 2017). This overfitting becomes notably evident in two-team zero-sum settings where both cooperative and competitive dynamics intertwine. A primary manifestation of the overfitting between cooperative agents, especially when all agents in the group share the same set of network parameters (i.e., parameter sharing (Foerster et al., 2016)), is in the agents becoming too accustomed to their training environments, leading to detailed coordination tailored to these specific conditions. As a consequence, when introduced to unfamiliar settings, their performance tends to falter. Concurrently, there is also an overfitting to specific opponent teams they have trained against. Instead of developing a flexible strategy that can withstand a variety of opponents, their strategies might be overly optimised to counteract the strategies of familiar adversaries. These dual forms of overfitting—both to the environment and to opponents—render such settings as perfect platforms to probe for vulnerabilities Tuyls et al. (2021). Furthermore, it is crucial to pinpoint a diverse set of adversarial scenarios for a holistic diagnostic of robustness, shedding light on possible shortcomings from various perspectives.
@@
Given these challenges, we introduce Multi-Agent Diagnostics for Robustness via Illuminated Diversity (MADRID), a novel method for systematically generating a diverse collection of adversarial settings where pre-trained multi-agent policies make strategic mistakes. To this end, MADRID employs approaches from quality-diversity (QD) (Lehman and Stanley, 2011; Cully and Demiris, 2018), a family of evolutionary algorithms that aim to generate a large collection of high-performing solutions each with their own unique characteristics.
@@
MADRID incorporates MAP-Elites (Mouret and Clune, 2015), a simple and effective QD approach, to systematically explore the vast space of adversarial settings. By discretising the search space, MADRID iteratively performs selection, mutation, and evaluation steps, continually refining and expanding the repertoire of high-performing adversarial scenarios within its archive (see Figure 1). A crucial feature of MADRID is its employment of the target policy’s regret— the gap in performance between the optimal and target policy—to quantify the quality of adversarial settings. Regret is shown to be an effective metric for identifying situations where RL agents underperform in both single-agent (Dennis et al., 2020; Jiang et al., 2021; Parker-Holder et al., 2022; Mediratta et al., 2023) and multi-agent (Samvelyan et al., 2023) domains. MADRID estimates a lower bound on the true regret by utilising a collection of reference policies Vinyals et al. (2019); Garnelo et al. (2021), which are not necessarily required to be high-performing. MADRID identifies situations where these reference policies surpass the target one, thereby providing a clear illustration of superior performance in given situations.
@@
To evaluate MADRID, we concentrate specifically on one of the most challenging multi-agent domains, namely the fully decentralised 11 vs 11 variation of Google Research Football (GRF, Kurach et al., 2020). This simulated environment is based on the popular real-world sport of football (a.k.a. soccer) and requires two teams of agents to combine short-term control techniques with coordinated long-term global strategies. GRF represents a unique combination of characteristics not present in other RL environments (Lin et al., 2023), namely multi-agent cooperation (within each team), competition (between the two teams), sparse rewards, large action and observation spaces, and stochastic dynamics. While many of the individual challenges in GRF, including multi-agent coordination (Rashid et al., 2018; Yu et al., 2022), long-term planning (Ecoffet et al., 2020) and non-transitivity (Balduzzi et al., 2019; Czarnecki et al., 2020), have been studied extensively in isolation, learning highly-competitive GRF policies has long remained outside the reach of RL methods. TiZero (Lin et al., 2023), a recent multi-agent RL approach, learned to ”master” the fully decentralised variation of GRF from scratch for the first time, using a hand-crafted curriculum, reward shaping, and self-play. Experimentally, TiZero has shown impressive results and outperformed previous methods by a large margin after an expensive training lasting 45 days on a large-scale distributed training infrastructure.
@@
We apply MADRID on GRF by targeting TiZero to diagnose a broad set of scenarios in which it commits tactical mistakes. Our extensive evaluations reveal diverse settings where TiZero exhibits poor performance, where weaker policies can outperform it. Specifically, MADRID discovers instances where TiZero is ineffective near the opponent’s goal, demonstrates a marked inability to comprehend the offside rule effectively, and even encounters situations of scoring accidental own goals. These findings highlight the latent vulnerabilities within even highly trained models and demonstrate that there is much room for improving their robustness. Our analysis showcases the value of identifying such adversarial settings in offering new insights into the hidden weaknesses of pretrained policies that may otherwise appear highly robust.
@@ 2. Background
In this work, we consider Underspecified Stochastic Games (USG), i.e., stochastic games (Shapley, 1953) with underspecified parameters of the environment. A USG for an environment with $n$ agents is defined by a set of states $\mathcal{S}$ and sets of per-agent actions $\mathcal{A}_{1},...,\mathcal{A}_{n}$ and observations $\mathcal{O}_{1},...,\mathcal{O}_{n}$. Each agent $i$ select actions using a stochastic policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\mapsto[0,1]$. $\Theta$ defines the set of free parameters of the environment that is incorporated into the transition function $\mathcal{T}:\mathcal{S}\times\Theta\times\mathcal{A}_{1}\times...\times\mathcal{A}_{n}\mapsto\mathcal{S}$ producing the next state based on the actions of all agents. Each agent $i$ receives observations $\mathbf{o}_{i}:\mathcal{S}\mapsto\mathcal{O}_{i}$ correlated with the current state and reward $r_{i}:\mathcal{S}\times\mathcal{A}_{i}\mapsto\mathbb{R}$ as a function of the state and agent’s action. The goal of each agent $i$ is to maximise its own total expected return $R_{i}=\sum_{t=0}^{T}\gamma^{t}r^{t}_{i}$ for the time horizon $T$, where $\gamma$ is a discount factor.
Each configuration of the free parameter $\theta\in\Theta$, which is often called a level (Jiang et al., 2021; Parker-Holder et al., 2022), defines a specific instantiation of the environment $\mathcal{M}_{\theta}$. For example, this can correspond to different positions of the walls in a maze, or locations of players and the ball in a football game. USG is a multi-agent variation of Underspecified POMDPs (Dennis et al., 2020) and the fully observable variant of UPOSGs (Samvelyan et al., 2023).
@@
Quality-diversity (QD) is a family of methods used to find a diverse collection of solutions that are performant and span a meaningful spectrum of solution characteristics (Lehman and Stanley, 2011; Cully and Demiris, 2018). The performance of solution $x\in\mathcal{X}$ is measure using the fitness $:\mathcal{X}\mapsto\mathbb{R}$ function. The diversity of solutions is typically measured using the feature_descriptor $:\mathcal{X}\mapsto\mathbb{B}$ function that maps a solution into the feature space $\mathbb{B}=\mathbb{R}^{K}$ that describes specific characteristics of the solution, such as behavioural properties or visual appearance.
Initialise: Empty $N$-dimensional grids for solutions $X$ and performances $\mathcal{P}$
Populate $n$ cells of $X$ with random solutions and corresponding cells of $P$ with their fitness scores
@@
MAP-Elites is a simple and effective QD method (Mouret and Clune, 2015). Here, the descriptor space $\mathbb{B}$ is discretised and represented as an initially empty $N\leq K$ dimensional grid, also referred to as the archive. The algorithm starts by generating an arbitrary collection of candidate solutions. At each iteration, a solution is randomly selected among those in the grid. A new solution is obtained by mutating the selected solution, which is then evaluated and mapped to a cell of the grid based on its feature descriptor. The solution is then placed in the corresponding cell of the grid if it has a higher fitness than the current occupant, or if the cell is empty. This cycle of selection, mutation, and evaluation is repeated, progressively enhancing both the diversity (coverage) and the quality (fitness) of the collection. The pseudo-code of MAP-Elites is presented in Algorithm 1.
@@ 3. MADRID
In this section, we describe Multi-Agent Diagnostics for Robustness via Illuminated Diversity (MADRID), a novel method for automatically generating diverse adversarial settings for a target pre-trained policy $\pi_{T}$. These are settings that either deceive the policy, forcing it to produce unintended behaviour, or where the policy inherently performs poorly, deviating from the optimal behaviour. For USGs, these settings correspond to particular environment levels $\theta\in\Theta$ that have been procedurally generated.
For quantifying adversarial levels, we make use of the target policy’s regret in level $\theta$, i.e., the difference in utility between the optimal policy $\pi^{*}$ and $\pi_{T}$ :
@@
where $V_{\theta}(\pi_{A},\pi_{B})=\mathop{\mathbb{E}}[\sum_{t=0}^{T}\gamma^{t}r_{t}^{A}]$ is the value of a policy $\pi_{A}$ against policy $\pi_{B}$ in $\theta$ [^2] Note that here, for the simplicity of the notation, we assume a two-team zero-sum setting. $\pi_{A}$ and $\pi_{B}$ describe the policies for groups of agents, either through a centralised controller or decentralised policies that employ parameter sharing. However, MADRID can be applied for more general multi-agent settings.
Regret is a suitable metric for evaluating adversarial examples in pre-trained models. It provides a measure that directly quantifies the suboptimality of a model’s decisions. While a high regret value serves as a glaring indicator of how far off a model’s behaviour is from the optimal choice, a low regret indicates the model’s decisions are closely aligned with the optimal choice. The importance of regret becomes even more pronounced when considering the varied scenarios in which a model might be deployed. Therefore, by investigating regret across diverse situations, we can not only pinpoint specific vulnerabilities of a model but also ensure robustness in previously unseen scenarios.
@@
Input: Target policy $\pi_{T}$, a collection of reference policies $\Pi_{R}$, $level\_descriptor:\Theta\mapsto\mathbb{R}^{N}$ function
@@
Initialise: Empty $N+1$-dimensional grids for levels $X$ and regret estimates $\mathcal{P}$
Populate $n$ cells of $X$ with randomly generated levels and corresponding estimated regret in $\mathcal{P}$
@@
MADRID casts the task of generating a diverse array of adversarial levels for each reference policy as a QD search problem. Specifically, MADRID uses MAP-Elites to systematically generate levels from $\Theta$ by discretising the feature space of levels into an $N$-dimensional grid, with an additional dimension representing the corresponding reference policy from $\Pi_{R}$. Using a discretised grid of MAP-Elites provides interpretability to the adversarial examples found in MADRID given that each cell defines specific environment parameters, alongside a reference policy which outperforms the target under these parameters.
MADRID starts by populating the grid with randomly generated initial levels for each reference policy. During the iterative process, levels are selected from the grid to undergo mutation, followed by regret estimation. Each mutated level is then mapped to a specific cell in the grid based on its features and replaces the existing occupant if the mutated level has higher regret or the corresponding cell is unoccupied. This procedure ensures a thorough exploration and exploitation of the environment design space, allowing MADRID to generate levels that are both diverse and have high regret. Figure 1 illustrates this process. Algorithm 2 provides the pseudocode of the method.
@@ 4. Experimental Setting
We evaluate MADRID on Google Research Football domain (GRF, Kurach et al., 2020). Our experiments seek to (1) showcase the effectiveness of MADRID in generating diverse adversarial settings for a target state-of-the-art pre-trained RL model, (2) analyse the adversarial settings generated by MADRID to find key weaknesses of the target model, (3) validate the design choices of MADRID by comparing it to two ablated baselines. Given its strong performance and usage in related works, Covariance Matrix Adaptation MAP-Elites (CMA-ME, Fontaine et al., 2020) serves as the base MAP-Elites method in our experiments. We provide full environment descriptions in Appendix B and implementation details in Appendix C.
@@
We compare MADRID against two baselines: The targeted baseline uses a MAP-Elites archive but randomly samples levels from scratch, rather than evolving previously discovered high-regret levels from the grid. Consequently, it does not leverage the discovered stepping stones throughout the search process (Lehman and Stanley, 2011). The random baseline samples levels randomly from scratch without maintaining an archive of high-regret levels.
@@
GRF is a complex open-source RL environment designed for training and evaluating agents to master the intricate dynamics of football, one of the world’s most celebrated sports. It offers a physics-based 3D simulation that tasks the RL policy with controlling a team of players to penetrate the opponent’s defence while passing the ball among teammates to score goals. GRF is a two-team zero-sum environment that has long been considered one of the most complex multi-agent RL benchmarks due to a unique combination of challenges (Huang et al., 2021; Wen et al., 2022; Lin et al., 2023), such as multi-agent cooperation, multi-agent competition, sparse rewards, large action and observation spaces, and stochastic dynamics. [^3] Highlighting the stochasticity of the GRF environment, a shot from the top of the box can lead to various outcomes, underscoring that not every action results in a predictable outcome.
In this work, we focus on the fully decentralised 11 vs 11 version of the environment where each of the $10$ RL agents on both sides controls an individual player on the field. [^4] The goalkeepers are controlled by the game AI. Following (Lin et al., 2023), each agent receives a $268$-dimensional feature vector as an observation including the player’s own characteristics, information about the ball, players of the both sides, as well as general match details. The action space of agents consists of $19$ discrete actions, such as moving in 8 directions, sprinting, passing, and shooting.
To apply MADRID on GRF, we utilise procedurally generated levels each represented as a vector consisting of $(x,y)$ coordinates of $20$ players [^5] The goalkeepers’ position positions are always near their own goals. and the ball. The position of the ball on the field serves as a convenient descriptor for levels in GRF because it accommodates diverse scenarios, ranging from attacking to defending on either side of the field. Therefore, we use the $x$ and $y$ coordinates of the ball as the two environment features in MADRID’s archive of levels. This leads to a categorisation of levels into $160$ uniformly spaced cells across the football field, as illustrated in Figure 3. Given that we are interested in evaluating TiZero in specific adversarial levels, in our experiments we restrict the episode length to $128$ steps taking place at the beginning of the game.
The third axis for the MAP-Elites archive indexes the reference policies $\Pi_{R}$. In our experiments, we make use of $48$ checkpoints of TiZero saved throughout its training (Lin et al., 2023), as well as three built-in bots in GRF with varying difficulties (easy, medium, and hard). For each reference policy, we initialise the grid with randomly sampled levels that assign random locations to players and the ball. Figure 3 illustrates some of the randomly generated levels.
@@
which corresponds to the difference in cross-play and self-play values between the reference and target policies. The performance on a given level $\theta$ between two policies $\pi_{A}$ and $\pi_{B}$ is the reward for scoring a goal:
@@
Figure 4. The estimated regret and goal score rate against TiZero in GRF are illustrated through each iteration for 51 reference agents, as shown in (a) and (c). The final values are presented in figures (b) and (d). Standard error over 3 random seeds is shown.